OpenStudy (anonymous):

The small pipe has a radius of 10 feet and a height of 22 feet while the large pipe has a height of 33 feet. What is the volume of the large pipe in terms of π?

OpenStudy (anonymous):

2,7475pi 2,200pi 4,500pi 7,425pi

OpenStudy (souvik):

is the radius of the large pipe is same with that of the small pipe?

OpenStudy (anonymous):

its not known

OpenStudy (souvik):

lets take its same... volume=pi*r^2*h....... here r= radius and h=height

OpenStudy (anonymous):

i keep using the formula and i looked up one online but i get the volume wrong

OpenStudy (souvik):

the radius is different then..

OpenStudy (anonymous):

ill retry

OpenStudy (jdoe0001):

if you could take a picture or screenshot of the material and post it here, that'd help

OpenStudy (anonymous):

they are similar so i set a proportions so i set them up n fond the LARGE radius is 15 plugged it in and did not get any of the multiple choice

OpenStudy (anonymous):

the choices are 2,7475pi 2,200pi 4,500pi 7,425pi

OpenStudy (jhannybean):



OpenStudy (anonymous):

its not inside of it

OpenStudy (jhannybean):

Create a ratio. \(V_1=V_2\)

OpenStudy (anonymous):

there are 2 diff

OpenStudy (jhannybean):

OH,wasn't specified. Ok.

OpenStudy (jdoe0001):

so give us a picture :/

OpenStudy (anonymous):

ok w8

OpenStudy (jhannybean):



OpenStudy (jhannybean):

\[\large \cfrac{33}{x} = \cfrac{22}{10}\]

OpenStudy (jhannybean):

\[\large x = 15\]

OpenStudy (jhannybean):

\[\large V_2 = \pi r^2 h\]\[\large V_2 = \pi (15)^2 (33)\]

OpenStudy (anonymous):

thats it

OpenStudy (jhannybean):

what is 225 x 33?

OpenStudy (jhannybean):

then just multiply by pi.

OpenStudy (anonymous):

7425

OpenStudy (jhannybean):

Good job.

OpenStudy (anonymous):

i set it up like you did but i got it wrong

OpenStudy (jhannybean):

Wait, it's not 7425?

OpenStudy (anonymous):

it is

OpenStudy (anonymous):

ive jdoe a medal,

OpenStudy (jhannybean):

You set it up like a proportion since it gives you the height and radius of 1 pipe, and the height of the other pipe but the missing radius. You find the length of the radius and use the formula for volume of a cylinder to calculate the volume of the bigger pipe.

OpenStudy (anonymous):

thank you

OpenStudy (nurali):

assuming the pipes are similar radius of large pipe = 10 * (33/22) = 15 volume of large pipe = pi *15^2 * 33 23314.5

OpenStudy (jhannybean):

O_o

OpenStudy (anonymous):

thats exactly what i got but turns out your not supposed to multiply pi so it was 7,425pi

OpenStudy (jhannybean):

Yep, in the problem they want it in terms of pi.

OpenStudy (nurali):

23314.5 or 7,425pi

OpenStudy (anonymous):

thanks guys

OpenStudy (jhannybean):

Np :)